OpenStudy (anonymous):

*help requested* *fan and medal awarded* Simplify 3x + 9 over 3 3x + 3 x + 3 x + 9 x What is the equation of the line that passes through the points (-2, 3) and (2, 7)? x - y = -1 x - y = -2 x - y = - 5 x - y = - 6 Evaluate: 18 + 4 ÷ 2 - 8 2 8 12 3

OpenStudy (anonymous):

ummm the last one is 3, im still trying to find out the others

OpenStudy (anonymous):

Divide each term separately\[\frac{ 3x + 9 }{ 3 }\]-----\[\frac{ 3x }{ 3 } = x\]\[\frac{ 9 }{ 3 } = 3\] x + 3

OpenStudy (anonymous):

2) Point-Slope form. y - y1 = m(x - x1) Find slope. (-2, 3) and (2, 7) \[m =\frac{ y2 - y1 }{ x2 - x1 }\]\[m = \frac{ 7 - 3 }{ 2 -(-2) } = \frac{ 7 - 3 }{ 2 + 2 } = \frac{ 4 }{ 4 } = 1\]Slope(m): 1 Point: (-2, 3) y - 3 = 1(x - (-2)) y - 3 = 1(x + 2) Change to Slope-intercept form. y = mx + b Distribute. y - 3 = x + 2 Add 3. y = x + 5 Change to standard form. Ax + By = C 'A' cannot be a negative or a fraction. y = x + 5 Subtract x from both sides. -x + y = 5 Multiply everything by -1. \[-1(-x + y = 5)\] x - y = -5 <-- Your answer. (:

OpenStudy (anonymous):

3) Use the order of operations: PEMDAS. - Parenthesis - Exponents - Multiplication/ Division - Addition/ Subtraction 18 + 4 ÷ 2 - 8 Divide 4 by 2. 18 + (4 ÷ 2) - 8 18 + 2 - 8 Add 18 and 2. 20 - 8 Subtract and Solve (: